### v3

# Credal Learning Theory

Abstract
Statistical learning theory is the foundation of machine learning, providing theoretical bounds for the risk of models learnt from a (single) training set, assumed to issue from an unknown probability distribution. In actual deployment, however, the data distribution may (and often does) vary, causing domain adaptation/generalization issues. In this paper we lay the foundations for a ‘credal’ theory of learning, using convex sets of probabilities (credal sets) to model the variability in the data-generating distribution. Such credal sets, we argue, may be inferred from a finite sample of training sets. Bounds are derived for the case of finite hypotheses spaces (both assuming realizability or not) as well as infinite model spaces, which directly generalize classical results.
1 Introduction
Statistical Learning Theory (SLT) considers the problem of predicting an output $y\in\mathcal{Y}$ given an input $x\in\mathcal{X}$ by means of a mapping $h:\mathcal{X}\rightarrow\mathcal{Y}$, $h\in\mathcal{H}$, called model or hypothesis, belonging to model (or hypotheses) space $\mathcal{H}$. The loss function $l:(\mathcal{X}\times\mathcal{Y})\times\mathcal{H}\rightarrow\mathbb{R}$ measures the error committed by a model $h\in\mathcal{H}$. For instance, the zero-one loss is defined as $l((x,y),h)\doteq\mathbbm{I}[y\neq h(x)]$, where $\mathbbm{I}$ denotes the indicator function. This function assigns a zero value to correct predictions and one to incorrect ones. Input-output pairs are usually assumed to be generated i.i.d. by a probability distribution $P^{\star}$, which is unknown. The expected risk – or expected loss – of the model $h$,
$L(h)\equiv L_{P^{\star}}(h)$ $\doteq\mathbb{E}_{P^{\star}}[l((x,y),h)]$
$=\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\star}(\text{d}(x,y)),$
measures the expected value – taken with respect to $P^{\star}$ – of loss $l$. The expected risk minimizer
$h^{\star}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}L(h),$
is any hypothesis in the given model space $\mathcal{H}$ that minimizes the expected risk. Given a training dataset $D=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ whose elements are drawn independent and identically distributed (i.i.d.) from probability distribution $P^{\star}$, the empirical risk of a hypothesis $h$ is the average loss over $D$. The empirical risk minimizer (ERM), i.e., the model $\hat{h}$ one actually learns from the training set $D$, is the one minimizing the empirical risk. Statistical Learning Theory seeks upper bounds on the expected risk $L(\hat{h})$ of the ERM $\hat{h}$, and in turn, on the excess risk, that is, on the difference between $L(\hat{h})$, and the lowest expected risk $L(h^{\star})$. This endeavor is pursued under increasingly more relaxed assumptions about the nature of the hypotheses space $\mathcal{H}$. Two common such assumptions are that either the model space is finite, or that there exists a model with zero expected risk (realizability).
In real-world situations, however, the data distribution may (and often does) vary, causing issues of domain adaptation (DA) or generalization (DG). Domain adaptation and generalization are interrelated yet distinctive concepts in machine learning, as they both deal with the challenges of transferring knowledge across different domains. The main goal of DA is to adapt a machine learning model trained on source domains to perform well on target domains. In opposition, DG aims to train a model that can generalize well to unseen data/domains not available during training. In simple terms, DA works on the assumption that our source and target domains are related to each other, meaning that they somehow follow a similar data generating probability distribution. DG, instead, assumes that the trained model should be able to handle unseen target data (Piva et al., 2023).
Some attempts to derive generalization bounds under more realistic conditions within classical SLT have been made. However, those approaches are characterized by a lack of generalizability, and the use of strong assumptions. A more detailed account of the state of the art and their limitations is discussed Section 2.
Figure 1: Graphical representation of the proposed learning framework. Given an available finite sample of training sets, each assumed to be generated by a single data distribution, one can learn a credal set $\mathcal{P}$ of data distributions in either a frequentist or subjectivist fashion (Section 3). This allows us to derive generalization bounds under credal uncertainty (Section 4).
Imprecise Probabilities (IPs) can provide a radically different solution to the construction of bounds in learning theory. A hierarchy of formalisms aimed at mathematically modeling the ‘epistemic’ uncertainty induced by sources such as lack of data, missing data or data which is imprecise in nature (Cuzzolin, 2020; Sale et al., 2023a, b), IPs have been successfully employed in the design of neural networks providing both better accuracy and uncertainty quantification to predictions (Sensoy et al., 2018; Manchingal and Cuzzolin, 2022; Manchingal et al., 2023; Caprio et al., 2023a; Lu et al., 2024; Wang et al., 2024). To date, however, they have never been considered as a tool to address the foundational issues of statistical learning theory associated with data drifting.
1.1 Contributions
This paper provides two innovative contributions: (1) the formal definition of a new learning setting in which models are inferred from a (finite) sample of training sets, rather than a single training set, each assumed to have been generated by a single data distribution (as in classical SLT); (2) the derivation of generalization bounds to the expected risk of a model learnt in this new learning setting, under the assumption that the epistemic uncertainty induced by the available training sets can be described by a credal set, i.e., a convex set of (data generating) probability distributions. The overall framework is illustrated in Figure 1.
Generalized upper bounds under credal uncertainty are derived under three increasingly realistic sets of assumptions, mirroring classical statistical learning theory treatment: (i) finite hypotheses spaces with realizability, (ii) finite hypotheses spaces without realizability, and (iii) infinite hypotheses spaces. We show that the corresponding classical results in SLT are special cases of the ones derived in the present paper.
1.2 Structure of the paper
The paper is structured as follows. First (Section 2) we present the existing work in addressing data distribution shifts in learning theory. We then introduce our new learning framework (Section 3). In Section 4 we illustrate the bounds derived under credal uncertainty, and show how classical results can be recovered as special cases. Section 5 concludes and outlines future undertakings. We prove our results in Appendix A, and we provide a synthetic experiment in Appendix B.
2 Related Work
The standard statistical approach is based on the assumption that test and training data are i.i.d. according to an unknown distribution. This assumption can fall short in real-world applications: as a result, many recent papers have been focusing on the “Out of Distribution” (OOD) generalization problem, also known as domain generalization (DG), to address the discrepancy between test and training distribution(s) (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021). Extensive surveys of existing methods and approaches to DG can be found in Zhou et al. (2022); Wang et al. (2022).
Although several proposals for learning bounds with theoretical guarantees have been made within DA, only a few attempts have been made in the field of DG (Redko et al., 2020; Deshmukh et al., 2019). Most of the theoretical attempts have focused on kernel methods, starting from the seminal work of Blanchard et al. (2011), spanning to a body of later work (see for example Muandet et al. (2013); Deshmukh et al. (2019); Hu et al. (2020)), or are reliant on strong assumptions (e.g. in relation to the similarity of training and testing domains) that are not directly applicable to broader DG approaches. Other work has focused on providing theoretical grounds showing that minimizing the maximum pairwise distance in the source domains (assumed to form a convex hull) leads to a distance decrease between any pair of domains in the convex hull (Albuquerque et al., 2019).
Researchers have also focused on adaptation to new domains over time, treating DG as an online game and the model as a player minimizing the risk associated with introducing new distributions by an adversary at each step (Rosenfeld et al., 2022). Ye et al. (2021) have attempted to relax assumptions and provide more general bounds, focusing on feature distribution characteristics; the authors have introduced terms related to stability and discriminative power to calculate, by an expansion function, the error bound on unseen domains. Causality principles have been leveraged in this sense, for example by Sheth et al. (2022) and Bellot and Bareinboim (2022), which provide distributional robustness guarantees using causal diagrams and source domain data. Researchers have also explored generalization bounds for DG based on the Rademacher complexity, allowing for the approach to be applicable to a broader range of models (Li et al., 2022).
On the empirical analysis side, Gulrajani and Lopez-Paz (2020) provide a comprehensive review of the state of the art. Though a simple ERM was found to outperform other more sophisticated methods (Cha et al., 2021), this approach is criticized for non-generalizability; Izmailov et al. (2018) highlights the importance of searching for flat minima in the training process for improved generalization.
All the aforementioned approaches take a deterministic, and point estimate-like, stance (i.e., assuming single training set) to the derivation of generalization bounds. In this paper, in opposition, we explicitly acknowledge the uncertainty inherent to domain variation in the form of a sample of training sets, each assumed to be generated by a different distribution, and propose a robust and flexible approach representing the resulting epistemic uncertainty via credal sets.
3 Credal Learning
Let us formalize the notion of learning a model from a collection of (training) sets of data, each issued from a different ‘domain’ characterized by a single, albeit unknown, data generating probability distribution.
Assume that we wish to learn a mapping (model, or hypothesis) $h:\mathcal{X}\rightarrow\mathcal{Y}$ between an input space $\mathcal{X}$ and an output space $\mathcal{Y}$ (where, once again, the mapping $h$ belongs to a hypotheses space $\mathcal{H}$), having as evidence a finite sample of training sets, $D_{1},\ldots,D_{N}$,
$D_{i}=\{(x_{i,1},y_{i,1}),\ldots,(x_{i,n_{i}},y_{i,n_{i}})\}.$
Assume also that the data in each set $D_{i}$ has been generated by a distinct probability distribution $P^{\star}_{i}$.
The question we want to answer is: What sort of guarantees can be derived on the expected risk of a model learnt from such a sample of training sets? How do they relate to classical Probably Approximately Correct (PAC) bounds from statistical learning theory?
3.1 Objectivist Modeling
While in classical statistical learning theory results are derived assuming no knowledge about the data generating process, the theorems and corollaries in this paper do require some knowledge, although incomplete, of the true distribution. To be more specific, we will posit that, by leveraging the available evidence $D_{1},\ldots,D_{N}$, the agent is able to elicit a credal set – i.e., a closed and convex set of probabilities – that contains the true data generating process $P^{\text{true}}$ for a new training sample $D_{N+1}$, possibly different from $D_{1},\ldots,D_{N}$. As we shall see in Section 4, though, this extra modeling effort allows us to derive stronger results.
There are at least two ways in which such credal set can be derived, that is, via objectivist and subjectivist modeling. In this section, we present the former, and we start by inspecting the frequentist approach to objectivist modeling.
3.1.1 Epsilon-contamination models
In classical frequentist statistics, given the available dataset, the agent assumes the analytical form of a likelihood $\mathcal{L}$ (not to be confused with the expected loss function, which we denote by a roman letter $L$), e.g. a Normal or a Gamma. As shown by Huber and Ronchetti (2009), though, small perturbations of the specified likelihood can induce substantial differences in the conclusions drawn from the data. A robust frequentist agent is interested in statistical methods that may not be fully optimal in the ideal model, but still lead to reasonable conclusions if the ideal model is only approximately true (Augustin et al., 2014). To account for this, the agent specifies the class of $\epsilon$-contaminated distributions $\mathcal{P}=\{P:P=(1-\epsilon)\mathcal{L}+\epsilon Q\text{, }\forall Q\}$, where $\epsilon$ is some positive quantity in $(0,1)$, and $Q$ is any distribution on $\mathcal{X}\times\mathcal{Y}$. Wasserman and Kadane (1990) show that $\mathcal{P}$ is indeed a (nonempty) credal set. In view of this robust frequentist goal, then, requiring that the true data generating process belongs to $\mathcal{P}$ is a natural assumption.
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, one approach to building the desired credal set is to specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$ and $\epsilon_{i}$-contaminate each of them to obtain $\mathscr{L}_{i}=\{P:P=(1-\epsilon_{i})\mathcal{L}_{i}+\epsilon_{i}Q\text{, }\forall Q\}$, $i\in\{1,\ldots,N\}$. The credal set $\mathcal{P}$ can then be derived by putting $\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i})$, where $\text{Conv}(\cdot)$ denotes the convex hull operator. [^1] It is easy to see that set $\mathcal{P}$ built this way is indeed a credal set. This because it is (i) convex by definition, and (ii) closed because it is the union of finitely many closed sets. An immediate consequence of Wasserman and Kadane (1990) and references therein is that $\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i})=\{P:P(A)\geq\underline{\mathcal{L}}(A)\text{, }\forall A\subseteq\mathcal{X}\times\mathcal{Y}\}$, where ¯$\underline{\mathcal{L}}(A)=\min_{i\in\{1,\ldots,N\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)$, for all $A\subseteq\mathcal{X}\times\mathcal{Y}$. A simple numerical example for this procedure is given in Appendix C.
We can also derive a credal set in an objectivist way by taking the plausibility function route. To describe it, we first need to introduce some concepts from the Dempster-Shafer theory of evidence (Dempster, 1967; Shafer, 1976).
3.1.2 Belief functions as lower probabilities
A random set (Kendall, 1974; Matheron, 1975; Nguyen, 1978; Molchanov, 2005) is a set-valued random variable, modeling random experiments in which observations come in the form of sets. In the case of finite sample spaces, they are called belief functions (Shafer, 1976). While classical discrete mass functions assign normalized, non-negative values to the elements $\omega\in\Omega$ of their sample space, a belief function independently assigns normalized, non-negative mass values to subsets of the sample space: $m(A)\geq 0$, for all $A\subseteq\Omega$, $\sum_{A\subseteq\Omega}m(A)=1$. The belief function associated with a mass function $m$ then measures the total mass of the subsets of each event $A$,
$\text{Bel}(A)=\sum_{B\subseteq A}m(B).\vspace{-3pt}$ (1)
Crucially, a belief function can be seen as the lower probability (or lower envelope) of the credal set
$\vspace{-3pt}\mathcal{M}(\text{Bel})=\Big{\{}P:\Omega\rightarrow[0,1]:\text{Bel}(A)\leq P(A)\text{, }\forall A\subseteq\Omega\Big{\}},\vspace{-0pt}$ (2)
where $P$ is a data distribution. The dual upper probability to Bel is $\text{Pl}(A)\doteq 1-\text{Bel}(A^{c})$, for all $A\subseteq\Omega$. When restricted to singleton elements, it is called contour function, $\text{pl}(\omega)=\text{Pl}(\{\omega\})$.
3.1.3 Inferring belief functions from data
There are various ways one can infer a belief (or, equivalently, a plausibility) function from (partial) data, such as a sample of training sets. If a classical likelihood $\mathcal{L}$ having probability density or mass function (pdf/pmf) $\ell$ is available (as assumed in the frequentist paradigm), [^2] Here, pdf/pmf $\ell$ is the result of the Radon-Nikodym derivative of $\mathcal{L}$ with respect to a sigma-finite dominating measure $\nu$. one can build a belief function by using the normalized likelihood as its contour function. That is,
$\text{pl}(\omega)\doteq\frac{\ell(\omega)}{sup_{\omega^{\prime}\in\Omega}\ell(\omega^{\prime})},\quad\forall\omega\in\Omega,$ (3)
where $\Omega=\mathcal{X}\times\mathcal{Y}$ is the space where the training pairs live.
As before, in our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we can specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$, and their corresponding pdf/pmf’s $\{\ell_{i}\}_{i=1}^{N}$. Then, we can compute ¯$\overline{\ell}(\omega)=\max_{i\in\{1,\ldots,N\}}\ell_{i}(\omega)$, for all $\omega\in\Omega$, and in turn $\text{pl}(\omega)=\overline{\ell}(\omega)/sup_{\omega^{\prime}\in\Omega}\overline{\ell}(\omega^{\prime})$, for all $\omega\in\Omega$ [^3] It is easy to see that pl is a well-defined plausibility contour. In turn, our credal set is derived as $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$, where $\text{d}P/\text{d}\nu=p$ is the pdf/pmf associated with distribution $P$ via its Radon-Nikodym derivative with respect to a sigma-finite dominating measure $\nu$. Such construction means that $\mathcal{P}$ includes all distributions whose pdf/pmf’s are element-wise dominated by plausibility contour pl.
3.1.4 Numerical example
Let us give a simple numerical example. Let $\Omega=\{\omega_{1},\omega_{2},\omega_{3}\}$, where $\omega_{j}=(x_{j},y_{j})$, $j\in\{1,2,3\}$. Suppose also that we observed four training samples $D_{1},\ldots,D_{4}$ and that we specified the likelihood pmf’s $\ell_{1},\ldots,\ell_{4}$ as in Table 1.444In this case, $\nu$ is the counting measure. It is immediate to see that ¯$\overline{\ell}=(0.4,0.8,0.6)^{\top}$. [^5] We write the upper likelihood $\overline{\ell}$ in vector form for notational convenience. Then, by (3), we have that $\text{pl}=(0.5,1,0.75)^{\top}$. We then derive the lower and upper probabilities, $\underline{P}$ and $\overline{P}$ respectively, of $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$ on $2^{\Omega}$ as in Table 2 (Augustin et al., 2014, Section 4.4). As we can see, the probability bounds imposed by the credal set are not too stringent, and in line with the evidence encapsulated in $\ell_{1},\ldots,\ell_{4}$. Hence, the assumption that $P^{\text{true}}\in\mathcal{P}$ is very plausible. For a visual representation of $\mathcal{P}$, see the yellow convex region in Figure 2.
Figure 2: Credal set generated as in the example.
3.1.5 Fiducial approach
An alternative approach, proposed by Dempster and Almond (Almond, 1992), is based on fiducial inference (Hannig, 2009). Consider a parametric model, i.e. a family of conditional probability distributions of the data
$\Big{\{}f(\omega|\theta):\omega\in\Omega,\theta\in\Theta\Big{\}},$ (4)
where $\Omega$ is, again, the observation space and $\Theta$ is a parameter space. If the parametric (sampling) model (4) is supplemented by a suitably designed auxiliary equation $\omega=a(\theta,u)$, where $u$ is a “pivot” variable of known a-priori distribution $\mu$, one obtains a random set
$\Gamma(u)=\{(\omega,\theta)\in\Omega\times\Theta:\omega=a(\theta,u)\}\subseteq\Omega\times\Theta$ (5)
which, in turn, induces a belief function on the product space $\Omega\times\Theta$. This can be finally be marginalized to the data space $\Omega=\mathcal{X}\times\mathcal{Y}$ to generate a lower probability on it.
This approach was further extended by Martin, Zhang and Liu, who used a “predictive” random set to express uncertainty on the pivot variable itself, leading to an inference technique called weak belief (Zhang and Liu, 2011).
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we can derive $N$ many random sets $\Gamma_{i}$ as in (5), $i\in\{1,\ldots,N\}$, and consider the $N$ belief functions $\text{Bel}_{i}$ on $\Omega\times\Theta$ they induce. Then, we can compute their marginalization $\text{Bel}_{i}|_{\Omega}$ on the data space $\Omega=\mathcal{X}\times\mathcal{Y}$, and compute the minimum $\underline{\text{Bel}}|_{\Omega}\doteq\min_{i\in\{1,\ldots,N\}}\text{Bel}_{i}|_{\Omega}$. [^6] It is easy to see that $\underline{\text{Bel}}|_{\Omega}$ is itself a well-defined belief function In turn, our credal set is given by $\mathcal{P}=\mathcal{M}(\underline{\text{Bel}}|_{\Omega})$ as in equation (2).
3.2 Subjectivist Modeling
Another way of specifying a credal set is by taking a personalistic (or subjectivist) route (Walley, 1991; Caprio and Mukherjee, 2021). In this instance, the agent proceeds as follows. First, they specify the lower probability $\underline{P}_{\mathcal{S}}$ on a finite collection $\mathcal{S}=\{A_{\mathcal{S}}\}$ of subsets of $\mathcal{X}\times\mathcal{Y}$ – i.e. the smallest value that the probability of set $A_{\mathcal{S}}$ can take on, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$. This can be done, for example, as a result of the empirical distribution.
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, let $\mathcal{S}=\cup_{i=1}^{N}D_{i}$, and $P^{\text{emp}}_{i}$ denote the empirical distribution associated with $D_{i}$. Before going on, let us add a notational remark. At the beginning of the section, we denoted by $P_{i}^{\star}$ the true data generating process for training set $D_{i}$, $i\in\{1,\ldots,N\}$. The empirical distribution $P^{\text{emp}}_{i}$ is a (non-parametric) estimation of $P_{i}^{\star}$. On the other hand, recall that we denoted by $P^{\text{true}}\equiv P^{\star}_{N+1}$ the true data generating process for training set $D_{N+1}$. Fix any $A_{\mathcal{S}}\in 2^{\mathcal{S}}$; its (non-vacuous) lower probability $\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$ is defined as
$\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})\doteq\min\left\{{P^{\text{emp}}_{i}(A_{\mathcal{S}}):P^{\text{emp}}_{i}(A_{\mathcal{S}})>0}\right\}.$
The lower probability $\underline{P}_{\mathcal{S}}$ assigns the least nonzero value derived from the $N$ available empirical distributions, to any subset $A_{\mathcal{S}}$ of the whole available evidence $\mathcal{S}=\cup_{i=1}^{N}D_{i}$.
3.2.1 Walley’s Natural Extension
Once a lower probability $\underline{P}_{\mathcal{S}}$ on $2^{\mathcal{S}}$ is inferred, it can be (coherently uniquely) extended to a lower probability $\underline{P}$ on the whole sigma-algebra endowed to $\mathcal{X}\times\mathcal{Y}$ through an operator called natural extension (Walley, 1991, Sections 3.1.7-3.1.9), (Troffaes and de Cooman, 2014). The resulting extended lower probability is such that $\underline{P}(A_{\mathcal{S}})=\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$, and a lower probability value $\underline{P}(A)$ is assigned to all the other subsets $A$ of $\mathcal{X}\times\mathcal{Y}$ that are not in $\mathcal{S}$. It is also coherent – in Walley’s terminology – because, in the behavioral interpretation of probability derived from de Finetti (1974, 1975), its values cannot be used to construct a bet that would make the agent lose for sure, no matter the outcome of the bet itself.
Once $\underline{P}$ is obtained, the agent can consider the core of $\underline{P}$,
$\mathcal{M}(\underline{P})\doteq\{P:\underline{P}(A)\leq P(A)\text{, }\forall A\subseteq\mathcal{X}\times\mathcal{Y}\},$
i.e., the collection of all the (countably additive) probabilities that set-wise dominate $\underline{P}$. Marinacci and Montrucchio (2004); Cerreia-Vioglio et al. (2015) have shown that $\mathcal{M}(\underline{P})$ is indeed a (nonempty) credal set.
3.2.2 Properties of the Core
As shown in Amarante and Maccheroni (2006, Example 1) and Amarante et al. (2006, Examples 6, 7, 8), given a generic credal set $\mathcal{Q}$ whose lower envelope is $\underline{Q}$ – i.e. a credal set $\mathcal{Q}$ for which $\underline{Q}(\cdot)=\inf_{Q\in\mathcal{Q}}Q(\cdot)$ – we have that $\mathcal{M}(\underline{Q})\supseteq\mathcal{Q}$. From an information-theoretic perspective, this means that the uncertainty encapsulated in the core of a lower probability $\underline{Q}$ is larger than that in any credal set whose lower envelope is $\underline{Q}$ (Gong and Meng, 2021; Caprio and Mukherjee, 2023; Caprio and Gong, 2023; Caprio and Seidenfeld, 2023; Caprio et al., 2023b). In turn, $\mathcal{M}(\underline{Q})$ is the largest credal set the agent can build which represents their partial knowledge.
In our learning framework, given the available evidence $D_{1},\ldots,D_{N}$ that the agent uses to derive $\underline{P}_{\mathcal{S}}$, if the agent is confident that $P^{\text{true}}(A_{\mathcal{S}})\geq\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$, then it is natural to assume $P^{\text{true}}\in\mathcal{M}(\underline{P})$.
4 Generalization Bounds under Credal Uncertainty
Consider a credal set $\mathcal{P}$ on $\mathcal{X}\times\mathcal{Y}$ derived as in Section 3, and assume that we collect new evidence in the form of training set $D_{N+1}=\{(x_{N+1,1},y_{N+1,1}),\ldots,(x_{N+1,n_{N+1}},y_{N+1,n_{N+1}})\}$. To ease notation, in the following we refer to the newly acquired evidence as $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$.
An assumption that is common to all the results we present in this section is that $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P\equiv P^{\text{true}}$ i.i.d., and $P\in\mathcal{P}$. This means that either the new evidence comes from one of the distributions that generated $D_{1},\ldots,D_{N}$, or that it is compatible with the credal set we built from past evidence (Gong and Meng, 2021). That is, either $P$ set-wise dominates the lower probability $\underline{P}$ of $\mathcal{P}$ like in sections 3.1, 3.1.2, and 3.2, or it is set-wise dominated by the upper probability $\overline{P}$ (induced e.g. by a contour function) of $\mathcal{P}$ like in section 3.1.3. This is an extremely natural assumption, especially when the flow of training sets that we collect pertains to similar experiments or tasks. We note in passing that assuming $P\in\mathcal{P}$ is less stringent than what is typically done in frequentist statistics, where the data generating process is assumed to be perfectly captured by the likelihood. We instead posit that the true data generating process for the new evidence available belongs to a credal set $\mathcal{P}$, that was derived by the sample of training sets $D_{1},\ldots,D_{N}$. We also point out how it could be checked whether the assumption $P\in\mathcal{P}$ holds, similarly e.g. to what Cella and Martin (2022, Section 7) and Javanmardi et al. (2024) do for credal-set-based conformal prediction methods. That being said, PAC-like guarantees on the correct distribution $P$ being an element of the credal set $\mathcal{P}$ need to be developed, a task that we defer to future work.
4.1 Realizability and Finite Hypotheses Space
Theorem 4.1.
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Let the empirical risk minimizer be
$\hat{h}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}l((x_{i},y_{i}),h).$ (6)
Assume that there exists a realizable hypothesis, that is, $h^{\star}\in\mathcal{H}$ such that $L_{P}(h^{\star})=0$, and that the model space $\mathcal{H}$ is finite. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Fix any $\delta\in(0,1)$. Then,
$\mathbb{P}\left[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)\right]\geq 1-\delta,$
where $\epsilon^{\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on extreme elements $\text{ex}\mathcal{P}$ of $\mathcal{P}$, i.e. those that cannot be written as a convex combination of one another.
Under some hypotheses, Theorem 4.1 gives us a tight probabilistic bound for the expected risk $L_{P}(\hat{h})$ of the empirical risk minimizer $\hat{h}$. [^7] Notice how, if $\mathcal{P}$ has finitely many extreme elements – which happens e.g. if we put $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{N})$, another frequentist way of deriving a credal set – then $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}$ is a finite union, and hence it is easier to compute. The bound holds for any possible distribution $P$ in the credal set $\mathcal{P}$ that generates the training data. A slightly looser bound depending on the diameter of credal set $\mathcal{P}$ holds if we calculate $L_{Q}(\hat{h})$ in place of $L_{P}(\hat{h})$.
Corollary 4.2.
Retain the assumptions of Theorem 4.1. Denote by $Q\in\mathcal{P}$, $Q\neq P$, a generic distribution in $\mathcal{P}$ different from $P$. Let $\Delta_{\mathcal{X}\times\mathcal{Y}}$ denote the space of all distributions over $\mathcal{X}\times\mathcal{Y}$, and endow it with the total variation metric $d_{TV}$. Then, pick any $\eta\in\mathbb{R}_{>0}$. If the diameter of $\mathcal{P}$, denoted by $\text{diam}_{TV}(\mathcal{P})$, is equal to $\eta$, we have that
$\mathbb{P}\left[L_{Q}(\hat{h})\leq\epsilon^{\star}(\delta)+\eta\right]\geq 1-\delta,$
where $\epsilon^{\star}(\delta)$ is the same quantity as in Theorem 4.1.
Corollary 4.2 gives us a probabilistic bound for the expected risk $L_{Q}(\hat{h})$ of the empirical risk minimizer $\hat{h}$, calculated with respect to the “wrong” distribution $Q$. That is, any distribution in $\mathcal{P}$ that is different from the one that generates the training set.
We now give a looser – but easier to compute – bound for $L_{P}(\hat{h})$.
Corollary 4.3.
Retain the assumptions of Theorem 4.1. We have that
$\epsilon^{\star}(\delta)\leq\epsilon_{\text{UB}}(\delta)\doteq\frac{\log|\mathcal{H}|+\log\left(\frac{1}{\delta}\right)}{n}.$
In turn,
$\mathbb{P}\left[L_{P}(\hat{h})\leq\epsilon_{\text{UB}}(\delta)\right]\geq 1-\delta,$
for all $P\in\Delta_{\mathcal{X}\times\mathcal{Y}}$.
Notice how $\epsilon_{\text{UB}}(\delta)$ is a uniform bound, that is, it holds for all possible distributions on $\mathcal{X}\times\mathcal{Y}$, not just those in $\mathcal{P}$. Strictly speaking, this means that we do not need to come up with a credal set $\mathcal{P}$ to find such a bound. Observe, though, that the bound $\epsilon^{\star}(\delta)$ that we derive in Theorem 4.1 is tighter because it is based on a procedure that leverages the evidence encoded in the credal set $\mathcal{P}$. A synthetic experiment confirming this, and studying other interesting properties of $\epsilon^{\star}(\delta)$ and $\epsilon_{\text{UB}}(\delta)$, can be found in Appendix B. By the proof of Theorem 4.1, we have that $L_{P}(\hat{h})$ behaves as $\mathcal{O}(\log|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|/n)$, which is faster than the rate $\mathcal{O}(\log|\mathcal{H}|/n)$ that we find in Corollary 4.3, since $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\subseteq\mathcal{H}$. Roughly, $B_{P^{\text{ex}}}$ is the set of “bad hypotheses” according to $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $L_{P^{\text{ex}}}(h)$ is larger than 00. A formal definition is given in the proof of Theorem 4.1. The modeling effort required by producing credal set $\mathcal{P}$ is rewarded with a tighter bound and a faster rate. Notice that Corollary 4.3 corresponds to Liang (2016, Theorem 4): we obtain a classical result as a special case of our more general theorem.
Let us now allow for distribution drift.
Corollary 4.4.
Consider a natural number $k<n$. Let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})\sim P_{1}$ i.i.d., and $(x_{k+1},y_{k+1}),\ldots,(x_{n},y_{n})\sim P_{2}$ i.i.d., where $P_{1},P_{2}$ are two generic elements of credal set $\mathcal{P}$. Retain the other assumptions of Theorem 4.1. Then,
$\mathbb{P}\left[L_{P_{1}}(\hat{h}_{1})+L_{P_{2}}(\hat{h}_{2})\leq\epsilon^{\star}(\delta)\frac{n^{2}}{k(n-k)}\right]\geq 1-\delta,$ (7)
where $\epsilon^{\star}(\delta)$ is the same quantity as in Theorem 4.1, and
$\hat{h}_{1}$ $\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\frac{1}{k}\sum_{i=1}^{k}l((x_{i},y_{i}),h),$
$\hat{h}_{2}$ $\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\frac{1}{n-k}\sum_{i=k+1}^{n}l((x_{i},y_{i}),h).$
Corollary 4.4 gives us a bound similar to the one in Theorem 4.1 when distribution drift is allowed. The price we pay for it is that it is looser. As a result of Corollary 4.3, for a looser but easier to compute bound, we can substitute $\epsilon^{\star}(\delta)$ with $\epsilon_{\text{UB}}(\delta)$.
4.2 No Realizability and Finite Hypotheses Space
Let us now relax the realizability assumption in Theorem 4.1.
Theorem 4.5.
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Assume that the model space $\mathcal{H}$ is finite. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Let also $\hat{h}$ be the empirical risk minimizer, and $h^{\star}$ be the best theoretical model. Fix any $\delta\in(0,1)$. Then,
$\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)\right]\geq 1-\delta,$
where $\epsilon^{\star\star}(\delta)$ is a well-defined quantity that depends only on $\delta$ and on the elements of $\text{ex}\mathcal{P}$.
We now show that also the “wrong” expected risk $L_{Q}(\hat{h})$ – that is, the expected risk computed according to $Q\in\mathcal{P}$ different from the one generating the training set – concentrates around the expected risk $L_{P}(h^{\star})$ evaluated at the best theoretical model $h^{\star}$.
Corollary 4.6.
Retain the assumptions of Theorem 4.5. Denote by $Q\in\mathcal{P}$, $Q\neq P$, a generic distribution in $\mathcal{P}$ different from $P$. Pick any $\eta\in\mathbb{R}_{>0}$; if the diameter $\text{diam}_{TV}(\mathcal{P})=\eta$, we have that
$\mathbb{P}\left[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)+\eta\right]\geq 1-\delta,$
where $\epsilon^{\star\star}(\delta)$ is the same quantity as in Theorem 4.5.
Similarly to Corollary 4.3, we can give a looser – but easier to compute – bound for $L_{P}(\hat{h})-L_{P}(h^{\star})$.
Corollary 4.7.
Retain the assumptions of Theorem 4.5. Then,
$\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)\doteq\sqrt{\frac{2\left(\log\left|\mathcal{H}\right|+\log\left(\frac{2}{\delta}\right)\right)}{n}}.$
In turn,
$\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\prime}_{\text{UB}}(\delta)\right]\geq 1-\delta,$
for all $P\in\Delta_{\mathcal{X}\times\mathcal{Y}}$.
The main difference with respect to Theorem 4.1 is that in Theorem 4.5, $L_{P}(\hat{h})-L_{P}(h^{\star})$ behaves as $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$, which is slower than what we had in Theorem 4.1. This is due to the relaxation of the realizability hypothesis. Just like before, though, we have that $\mathcal{O}(\sqrt{\log|B^{\prime}_{\text{ex}\mathcal{P}}|/n})$ is faster than the rate $\mathcal{O}(\sqrt{\log|\mathcal{H}|/n})$ that we find in Corollary 4.7. This because $B^{\prime}_{\text{ex}\mathcal{P}}\subseteq\mathcal{H}$. Roughly, $B^{\prime}_{\text{ex}\mathcal{P}}$ is the set of “bad hypotheses” according to at least one $P^{\text{ex}}\in\text{ex}\mathcal{P}$. That is, those $h$’s for which $|\hat{L}_{P^{\text{ex}}}(h)-L_{P^{\text{ex}}}(h)|$ is larger than 00, for at least one $P^{\text{ex}}$. A formal definition is given in the proof of Theorem 4.5. Notice that Corollary 4.7 corresponds to Liang (2016, Theorem 7): we obtain a classical result as a special case of our more general theorem.
Let us now allow for distribution drift. To improve notation clarity, in the following we let $h^{\star}_{P}$ denote an element of $\operatorname*{arg\,min}_{h\in\mathcal{H}}L_{P}(h)$, for a distribution $P\in\mathcal{P}$.
Corollary 4.8.
Consider a natural number $k<n$. Let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})\sim P_{1}$ i.i.d., and $(x_{k+1},y_{k+1}),\ldots,(x_{n},y_{n})\sim P_{2}$ i.i.d., where $P_{1},P_{2}$ are two generic elements of credal set $\mathcal{P}$. Retain the other assumptions of Theorem 4.5. Then,
$\mathbb{P}\bigg{[}\bigg{(}$ $L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}\left(h^{\star}_{P_{1}}\right)\bigg{)}+\bigg{(}L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}\left(h^{\star}_{P_{2}}\right)\bigg{)}$
$\leq\epsilon^{\star\star}(\delta)\sqrt{\frac{n}{k(n-k)}}\left(\sqrt{k}+\sqrt{n-k}\right)\bigg{]}\geq 1-\delta,$
where $\epsilon^{\star\star}(\delta)$ is the same quantity as in Theorem 4.5, and $\hat{h}_{1}$ and $\hat{h}_{2}$ are defined as in Corollary 4.4.
Corollary 4.8 tells us that the excess risk is bounded also in the presence of distribution drift. The price we pay for allowing distribution shift is a looser bound. As a result of Corollary 4.7, for a looser but easier to compute bound, we can substitute $\epsilon^{\star\star}(\delta)$ with $\epsilon^{\prime}_{\text{UB}}(\delta)$.
4.3 No Realizability and Infinite Hypotheses Space
We now relax also the finite hypotheses space assumption in Theorem 4.1.
Theorem 4.9.
Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P$ i.i.d., where $P$ is any element of credal set $\mathcal{P}$. Let $l$ denote the zero-one loss, i.e. $l((x,y),h)=\mathbbm{I}[y\neq h(x)]$. Let also $\hat{h}$ be the empirical risk minimizer, and $h^{\star}$ be the best theoretical model. Fix any $\delta\in(0,1)$. Then,
$\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star\star}(\delta)\right]\geq 1-\delta,$ (8)
for all $P\in\mathcal{P}$. Here,
$\epsilon^{\star\star\star}(\delta)\doteq 4\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{n}},$
where
$\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})\doteq\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$
and
$R_{n,P^{\text{ex}}}(\mathcal{A})\doteq\mathbb{E}_{P^{\text{ex}}}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}l((x_{i},y_{i}),h)\right].$ (9)
Here, $\sigma_{1},\ldots,\sigma_{n}\sim\text{Unif}(\{-1,1\})$, and $\mathcal{A}\doteq\{(x,y)\mapsto l((x,y),h):h\in\mathcal{H}\}$. [^8] Class $\mathcal{A}$ is the loss class, and it is the composition of the zero-one loss function with each of the hypotheses in $\mathcal{H}$ (Liang, 2016, Page 70).
Notice that $R_{n,P^{\text{ex}}}(\mathcal{A})$ is a slight modification of the classical definition of the Rademacher complexity of class $\mathcal{A}$ [^9] Recall that the Rademacher complexity of $\mathcal{A}$ measures how well the best element of $\mathcal{H}$ fits random noise (coming from the $\sigma_{i}$’s) (Liang, 2016, Page 69), (Balcan and Berlind, 2011). Indeed, the latter is given by
$R_{n}(\mathcal{A})\equiv R_{n,P}(\mathcal{A})\doteq\mathbb{E}_{P}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}l((x_{i},y_{i}),h)\right].$
This because the expectation is taken with respect to the same distribution $P$ according to which the training elements $(x_{1},y_{1}),\ldots,$ $(x_{n},y_{n})$ are drawn. We consider $R_{n,P^{\text{ex}}}(\mathcal{A})$ instead of $R_{n,P}(\mathcal{A})$ because, since $\mathcal{P}$ is a credal set, we have that $P$ can be written as a convex combination of the extreme elements of $\mathcal{P}$, and $\sup_{P\in\mathcal{P}}R_{n,P}(\mathcal{A})=\sup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}R_{n,P^{\text{ex}}}(\mathcal{A})$. If the credal set is finitely generated, that is, if it has finitely many extreme elements, [^10] See footnote 7. then it is easier to compute $\epsilon^{\star\star\star}(\delta)$: we only need to compute a maximum in place of a supremum.
As we show in Corollary 4.10, Theorem 4.9 generalizes Liang (2016, Theorem 9). This latter focuses only on the “true” probability $P^{\star}$ on $\mathcal{X}\times\mathcal{Y}$, while our result holds for all the plausible distributions in credal set $\mathcal{P}$. This grants us to hedge against distribution misspecification. Let us pause here to add a clarification. In real-world applications, we effectively cannot compute $R_{n,P^{\star}}(\mathcal{A})$, since distribution $P^{\star}$ is unknown. While $R_{n,P^{\star}}(\mathcal{A})$ can be approximated via the empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$ (Liang, 2016, Equation (219)), whose expected value is indeed $R_{n,P^{\star}}(\mathcal{A})$, doing so has at least two drawbacks:
(1) Especially in the case of low cardinality training set, i.e., if $n\equiv n_{D_{N+1}}$ is not “large enough”, this may lead to a poor approximation of the classical bound (equation (13) in Appendix A);
(2) The collected dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ may well be a realization from the tail of distribution $P^{\star}$. The empirical Rademacher complexity $\hat{R}_{n}(\mathcal{A})$, then, would be a poor approximation of $R_{n,P^{\star}}(\mathcal{A})$.
Instead, while $\overline{R}_{n,P^{\text{ex}}}(\mathcal{A})$ is more conservative, it can be computed explicitly – since we know credal set $\mathcal{P}$ and its extreme elements $\text{ex}\mathcal{P}$ – and it leads to a bound that, although looser, holds for all $P\in\mathcal{P}$.
Corollary 4.10.
Retain the assumptions of Theorem 4.9. If $\mathcal{P}$ is the singleton $\{P^{\star}\}$, we retrieve Liang (2016, Theorem 9).
We then derive a more general version of Corollary 4.6.
Corollary 4.11.
Retain the assumptions of Theorem 4.9. Denote by $Q\in\mathcal{P}$, $Q\neq P$, a generic distribution in $\mathcal{P}$ different from $P$. Pick any $\eta\in\mathbb{R}_{>0}$; if $\text{diam}_{TV}(\mathcal{P})=\eta$, we have that
$\mathbb{P}\left[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star\star}(\delta)+\eta\right]\geq 1-\delta,$
where $\epsilon^{\star\star\star}(\delta)$ is the same quantity as in Theorem 4.9.
Let us now allow for distribution drift.
Corollary 4.12.
Consider a natural number $k<n$. Let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})\sim P_{1}$ i.i.d., and $(x_{k+1},y_{k+1}),\ldots,(x_{n},y_{n})\sim P_{2}$ i.i.d., where $P_{1},P_{2}$ are two generic elements of credal set $\mathcal{P}$. Retain the other assumptions of Theorem 4.9, and let
$\epsilon^{\star\star\star}_{\text{shift}}$ $\doteq 4\left[\overline{R}_{k,P^{\text{ex}}}(\mathcal{A})+\overline{R}_{n-k,P^{\text{ex}}}(\mathcal{A})\right]$
$+\sqrt{\frac{2\log(2/\delta)}{n(n-k)}}\left(\sqrt{n-k}+\sqrt{n}\right).$
Then,
$\mathbb{P}\bigg{[}\bigg{(}$ $L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}\left(h^{\star}_{P_{1}}\right)\bigg{)}$
$+\left(L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}\left(h^{\star}_{P_{2}}\right)\right)\leq\epsilon^{\star\star\star}_{\text{shift}}\bigg{]}\geq 1-\delta,$
where $\hat{h}_{1}$ and $\hat{h}_{2}$ are defined as in Corollary 4.4.
Similar considerations as the ones after Corollaries 4.4 and 4.8 hold in this more general case as well.
5 Conclusions
In this paper, we laid the foundations of a more general statistical learning theory, that we called Credal Learning Theory (CLT). We generalized some of the most important results of classical Statistical Learning Theory (SLT) to allow for drift and misspecification of the data generating process. We did so by considering sets of probabilities (credal sets), instead of single distributions. The modeling effort needed to elicit credal sets is paid off in terms of the tightness of the resulting bounds. This is just the first step in an ongoing endeavor to develop a comprehensive apparatus that extends all the classical SLT results to a credal sets system. In the future, we plan to further our undertaking, for instance by (i) modeling the epistemic uncertainty induced by domain variation through random sets rather than credal sets, (ii) comparing our method with robust learning (Cianfarani et al., 2022), and (iii) extending our results to different losses. In this work, we focused solely on the zero-one loss to be able to directly build on the classical results in Liang (2016, Chapter 3). We also intend to validate empirically our findings. For now, a synthetic experiment showing that the bound in Theorem 4.1 is tighter than the classical SLT one, can be found in Appendix B.
Appendix A Proofs
Proof of Theorem 4.1.
[…]
$\mathbb{P}[\hat{h}\in B_{P}]\leq\mathbb{P}[\exists h\in B_{P}:\hat{L}({h})=0].$
Let us bound $\mathbb{P}[\hat{L}({h})=0]$ for a fixed $h\in B_{P}$. Given our choice of zero-one loss, on each example, hypothesis $h$ does not err with probability $1-L_{P}(h)$. Since the training examples are i.i.d. and $L_{P}(h)>\epsilon$ for all $h\in B_{P}$, then
$\mathbb{P}[\hat{L}({h})=0]=(1-L_{P}(h))^{n}\leq(1-\epsilon)^{n}\leq\exp(-\epsilon n).$ (10)
Applying the union bound, we obtain
$\mathbb{P}[\exists h\in B_{P}:\hat{L}({h})=0]$ $\leq\sum_{h\in B_{P}}\mathbb{P}[\hat{L}({h})=0]$
$\leq|B_{P}|\exp(-\epsilon n)$
$\leq|\cup_{P\in\mathcal{P}}B_{P}|\exp(-\epsilon n)$
$=|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|\exp(-\epsilon n)$
$\doteq\delta.$
The penultimate equality comes from $\mathcal{P}$ being a credal set, by the Bauer Maximum Principle and the linearity of the expectation operator. Rearranging the terms we get
$\epsilon^{\star}(\delta)\equiv\epsilon=\frac{\log|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|+\log(1/\delta)}{n}.$
In turn, this implies that $\mathbb{P}[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)]\geq 1-\delta$. ∎
Proof of Corollary 4.2.
Let $A_{\hat{h}}\doteq\{(x,y)\in\mathcal{X}\times\mathcal{Y}:y\neq\hat{h}(x)\}\in\mathcal{A}_{\mathcal{X}\times\mathcal{Y}}$. Notice that
$L_{P}(\hat{h})$ $=\mathbb{E}_{P}[\mathbb{I}(y\neq\hat{h}(x))]=P(A_{\hat{h}}),$
$L_{Q}(\hat{h})$ $=\mathbb{E}_{Q}[\mathbb{I}(y\neq\hat{h}(x))]=Q(A_{\hat{h}}).$
Recall that $\text{diam}_{TV}(\mathcal{P})\doteq\sup_{P,Q\in\mathcal{P}}\sup_{A\in\mathcal{A}_{\mathcal{X}\times\mathcal{Y}}}|P(A)-Q(A)|$. As a consequence, we have that $L_{Q}(\hat{h})=L_{P}(\hat{h})+\zeta_{Q}$, where $\zeta_{Q}$ is a quantity in $[-\eta,\eta]$ depending on $Q$. Given our assumption on the diameter, then, $L_{P}(\hat{h})+\zeta_{Q}\leq L_{P}(\hat{h})+\eta$, so $L_{Q}(\hat{h})-\eta\leq L_{P}(\hat{h})$. In turn this implies that
$\mathbb{P}\left[L_{Q}(\hat{h})-\eta\leq\epsilon^{\star}(\delta)\right]\geq\mathbb{P}\left[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)\right].$
The proof is concluded by noting that $\mathbb{P}[L_{Q}(\hat{h})-\eta\leq\epsilon^{\star}(\delta)]=\mathbb{P}[L_{Q}(\hat{h})\leq\epsilon^{\star}(\delta)+\eta]$, and that $\mathbb{P}[L_{P}(\hat{h})\leq\epsilon^{\star}(\delta)]\geq 1-\delta$ by Theorem 4.1. ∎
Proof of Corollary 4.3.
Since $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\subseteq\mathcal{H}$, it is immediate to see that $\epsilon^{\star}(\delta)\leq\epsilon_{\text{UB}}(\delta)$. In turn,
$\mathbb{P}\left[\sup_{P\in\mathcal{P}}L_{P}(\hat{h})\leq\epsilon_{\text{UB}}(\delta)\right]\geq 1-\delta,$
or equivalently, $\mathbb{P}[L_{P}(\hat{h})\leq\epsilon_{\text{UB}}(\delta)]\geq 1-\delta$, for all $P\in\mathcal{P}$. ∎
Proof of Corollary 4.4.
From Theorem 4.1, we have that
$\mathbb{P}\left[L_{P_{1}}(\hat{h}_{1})\leq\frac{\log\left|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\right|+\log(1/\delta)}{k}\right]\geq 1-\delta,$
and that
$\mathbb{P}\left[L_{P_{2}}(\hat{h}_{2})\leq\frac{\log\left|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}\right|+\log(1/\delta)}{n-k}\right]\geq 1-\delta.$
The result, then, is an immediate consequence of the additivity of the expectation operator and of probability $\mathbb{P}$. ∎
Proof of Theorem 4.5.
The proof builds on that of Liang [2016, Theorem 7]. Fix any $\epsilon>0$, and any $P\in\mathcal{P}$. Assume that the training dataset is given by $n$ i.i.d. draws from $P$. By Liang [2016, Equations (158) and (186)], we have that
$\begin{split}\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}&\leq\mathbb{P}\left[\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]\\ &<|\mathcal{H}|\cdot 2\exp\left(-2n\left(\frac{\epsilon}{2}\right)^{2}\right)\\ &\doteq\delta(\epsilon).\end{split}$ (11)
Notice though, that we can improve on this bound, since we know that $P\in\mathcal{P}$, a credal set. Let $B^{\prime}_{P}\doteq\{h\in\mathcal{H}:|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}$ be the set of “bad hypotheses” according to $P$. Then, it is immediate to see that
$\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{P}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|.$
Notice though that we do not know $P$; we only know it belongs to $\mathcal{P}$. Hence, we need to consider the set $B^{\prime}_{\mathcal{P}}$ of bad hypotheses according to all the elements of $\mathcal{P}$, that is, $B^{\prime}_{\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P\in\mathcal{P}\text{,}|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P\in\mathcal{P}}B^{\prime}_{P}$. Since $\mathcal{P}$ is a credal set, by the Bauer Maximum Principle and the linearity of the expectation operator we have that $B^{\prime}_{\mathcal{P}}=B^{\prime}_{\text{ex}\mathcal{P}}\doteq\{h\in\mathcal{H}:\exists P^{\text{ex}}\in\text{ex}\mathcal{P}\text{, }|\hat{L}_{P}({h})-L_{P}(h)|>\epsilon/2\}=\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B^{\prime}_{P^{\text{ex}}}$. Hence, we obtain
$\sup_{h\in\mathcal{H}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|=\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|.$
In turn, (11) implies that
$\mathbb{P}\big{[}L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon\big{]}$ $\leq\mathbb{P}\left[\sup_{h\in B^{\prime}_{\text{ex}\mathcal{P}}}\left|\hat{L}_{P}({h})-L_{P}(h)\right|>\frac{\epsilon}{2}\right]$
$<|B^{\prime}_{\text{ex}\mathcal{P}}|\cdot 2\exp\left(-2n\left(\frac{\epsilon}{2}\right)^{2}\right)$
$\doteq\delta_{\text{ex}\mathcal{P}}.$
Rearranging, we obtain
$\epsilon=\sqrt{\frac{2\left(\log\left|B^{\prime}_{\text{ex}\mathcal{P}}\right|+\log\left(\frac{2}{\delta_{\text{ex}\mathcal{P}}}\right)\right)}{n}},$ (12)
so if $\delta$ is fixed, we can write $\epsilon\equiv\epsilon^{\star\star}(\delta)$. In turn, this implies that $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})>\epsilon^{\star\star}(\delta)]<\delta$, or equivalently, $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)]\geq 1-\delta$. ∎
Proof of Corollary 4.6.
The first part of the proof is very similar to that of Corollary 4.2. Given our assumption on the diameter, we have that $L_{Q}(\hat{h})-L_{P}(h^{\star})=L_{P}(\hat{h})+\zeta_{Q}-L_{P}(h^{\star})$, where $\zeta_{Q}$ is a quantity in $[-\eta,\eta]$ depending on $Q$. Then, $L_{P}(\hat{h})+\zeta_{Q}-L_{P}(h^{\star})\leq L_{P}(\hat{h})+\eta-L_{P}(h^{\star})$, so $L_{Q}(\hat{h})-\eta-L_{P}(h^{\star})\leq L_{P}(\hat{h})-L_{P}(h^{\star})$. In turn this implies that $\mathbb{P}\left[L_{Q}(\hat{h})-\eta-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)\right]\geq\mathbb{P}\left[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)\right]$. The proof is concluded by noting that $\mathbb{P}[L_{Q}(\hat{h})-\eta-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)]=\mathbb{P}[L_{Q}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)+\eta]$, and that $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\star\star}(\delta)]\geq 1-\delta$ by Theorem 4.5. ∎
Proof of Corollary 4.7.
Since $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B^{\prime}_{P^{\text{ex}}}\subseteq\mathcal{H}$, it is immediate to see that $\epsilon^{\star\star}(\delta)\leq\epsilon^{\prime}_{\text{UB}}(\delta)$. In turn,
$\mathbb{P}\left[\sup_{P\in\mathcal{P}}\left(L_{P}(\hat{h})-L_{P}(h^{\star})\right)\leq\epsilon^{\prime}_{\text{UB}}(\delta)\right]\geq 1-\delta,$
or equivalently, $\mathbb{P}[L_{P}(\hat{h})-L_{P}(h^{\star})\leq\epsilon^{\prime}_{\text{UB}}(\delta)]\geq 1-\delta$, for all $P\in\mathcal{P}$. ∎
Proof of Corollary 4.8.
From Theorem 4.5, we have that
$\mathbb{P}\left[L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}\left(h^{\star}_{P_{1}}\right)\leq\sqrt{\frac{2\left(\log\left|B^{\prime}_{\text{ex}\mathcal{P}}\right|+\log\left(\frac{2}{\delta}\right)\right)}{k}}\right]\geq 1-\delta,$
and that
$\mathbb{P}\left[L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}\left(h^{\star}_{P_{2}}\right)\leq\sqrt{\frac{2\left(\log\left|B^{\prime}_{\text{ex}\mathcal{P}}\right|+\log\left(\frac{2}{\delta}\right)\right)}{n-k}}\right]\geq 1-\delta.$
The result, then, is an immediate consequence of the additivity of the expectation operator and of probability $\mathbb{P}$. ∎
Proof of Theorem 4.9.
Fix any $\delta\in(0,1)$. In Liang [2016, Theorem 9], the author shows that for a fixed probability measure $P$ on $\mathcal{X}\times\mathcal{Y}$, we have that
$L_{P}(\hat{h})-L_{P}(h^{\star})\leq 4R_{n,P}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{n}}$ (13)
holds with probability at least $1-\delta$, where $R_{n,P}$ is defined analogously as in (9). The result in (8), then, follows from $\mathcal{P}$ being a credal set, and the expectation being a linear operator. ∎
Proof of Corollary 4.10.
Immediate from Theorem 4.9. ∎
Proof of Corollary 4.11.
The proof is very similar to that of Corollary 4.6. ∎
Proof of Corollary 4.12.
From Theorem 4.9, we have that
$\mathbb{P}\left[L_{P_{1}}(\hat{h}_{1})-L_{P_{1}}(h^{\star}_{1})\leq 4\overline{R}_{k,P^{\text{ex}}}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{k}}\right]\geq 1-\delta,$
and that
$\mathbb{P}\left[L_{P_{2}}(\hat{h}_{2})-L_{P_{2}}(h^{\star}_{2})\leq 4\overline{R}_{n-k,P^{\text{ex}}}(\mathcal{A})+\sqrt{\frac{2\log(2/\delta)}{n-k}}\right]\geq 1-\delta.$
The result, then, is an immediate consequence of the additivity of the expectation operator and of probability $\mathbb{P}$. ∎
Appendix B A Synthetic Experiment on Theorem 4.1
In this section, we perform a synthetic experiment to show that the bound we find in Theorem 4.1 is indeed tighter than the classical SLT one reported in Corollary 4.3. In recent literature, studies by Kacham and Woodruff [2022], Amit et al. [2022], Li and Liu [2021] have conducted synthetic experiments in a similar manner. These works are mainly theoretical in nature, but they also acknowledge the importance of experimental validation with preliminary analysis.
Experiment 1: Let the available training sets be $D_{1},D_{2},D_{3}$. Assume, for simplicity, that $\Omega=\mathcal{X}\times\mathcal{Y}=\{x\}\times\mathbb{R}\simeq\mathbb{R}$. Suppose that we specified the likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$. Call $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ their respective probability measures, and derive the credal set $\mathcal{P}$ as we did in footnote 7. That is, let $\mathcal{P}=\text{Conv}(\{\mathcal{L}_{i}\}_{i=1}^{3})$. We determine the credal set in this way because it is then easy to find its extreme elements $\text{ex}\mathcal{P}$. Indeed, it is immediate to notice that $\text{ex}\mathcal{P}=\{\mathcal{L}_{i}\}_{i=1}^{3}$. Let now $D_{N+1}\equiv D_{4}$ be a collection of $n$ samples from $P^{\text{true}}\equiv\mathcal{L}_{2}\in\mathcal{P}$. The hypothesis space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and we consider the zero-one loss $l$ as we did in the main portion of the paper.
We need to find $\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}=\cup_{i=1}^{3}\{h\in\mathcal{H}:L_{\mathcal{L}_{i}}(h)>\epsilon\}$, where $\epsilon$ depends on $\delta$ as in the proof of Theorem 4.1. That is, we want those $h$’s for which the expected loss according to $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ or $\mathcal{L}_{3}$ is larger than $\epsilon$. They are the collection of “bad hypotheses” according to at least one of the extreme elements of our credal set. Recall that
$\epsilon^{\star}(\delta)=\frac{\log|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|+\log(1/\delta)}{n}$
is the bound we found in Theorem 4.1, and that
$\epsilon_{\text{UB}}(\delta)=\frac{\log|\mathcal{H}|+\log(1/\delta)}{n}$
is the classical SLT bound, that we reported in Corollary 4.3.
As we can see from Table 1, our bound $\epsilon^{\star}(\delta)$ improves on the classical SLT one $\epsilon_{\text{UB}}(\delta)$. Table 1 also tells us that bound $\epsilon^{\star}(\delta)$ is tighter than $\epsilon_{\text{UB}}(\delta)$ when the sample size $n=|D_{4}|$ is small, and then $\epsilon^{\star}(\delta)$ becomes progressively closer to $\epsilon_{\text{UB}}(\delta)$ as $n=|D_{4}|$ increases. This same pattern is observed when the extrema of the credal set are closer to each other. Indeed, in Table 2 we repeat the experiment and choose as extrema of $\mathcal{P}$ three measures whose pdf’s are three Normals $\mathcal{N}(-0.1,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(0.1,1)$. [^12] Of course, they are much closer to each other than the Normals $\mathcal{N}(-5,1)$, $\mathcal{N}(0,1)$, and $\mathcal{N}(5,1)$, e.g. in the Total Variation metric. The reason for this behavior is the following. With few available samples, that is, when $n=|D_{4}|$ is low, Credal Learning Theory is able to leverage the evidence encoded in the credal set, and hence to derive a tighter bound than classical Statistical Learning Theory. When the sample size is large, that is, when $n=|D_{4}|$ is high, the classical bound $\epsilon_{\text{UB}}(\delta)$ itself is very small. This because the amount of evidence available is large, and so $1/n\sum_{i=1}^{n}l((x_{i},y_{i}),h)$ well approximates $\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\text{true}}(\text{d}(x,y))$. In turn, since $\epsilon_{\text{UB}}(\delta)$ is already very small, then the CLT bound $\epsilon^{\star}(\delta)$ we derive cannot improve greatly on it. Hence, their values are close together, despite $\epsilon^{\star}(\delta)$ being slightly tighter. The code for this experiment is available upon request.
Table 1: Results of experimental evaluation of our bound tightness. Here the hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-5,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(5,1)$.
# Samples $n$ $\epsilon^{\star}(\delta)$ $\epsilon_{\text{UB}}(\delta)$ $|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|$ $|\mathcal{H}|$ Realizability
$10$ $0.74500$ $0.76009$ $86$ $100$ Yes
$100$ $0.07560$ $0.07600$ $96$ $100$ Yes
$200$ $0.03785$ $0.03800$ $97$ $100$ Yes
$300$ $0.02526$ $0.02533$ $98$ $100$ Yes
$400$ $0.01897$ $0.01900$ $99$ $100$ Yes
$500$ $0.01516$ $0.01520$ $98$ $100$ Yes
Table 2: Results of experimental evaluation of our bound tightness. Here the hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$. The likelihood pdfs $\ell_{1}=\mathcal{N}(-0.1,1)$, $\ell_{2}=\mathcal{N}(0,1)$, and $\ell_{3}=\mathcal{N}(0.1,1)$.
# Samples $n$ $\epsilon^{\star}(\delta)$ $\epsilon_{\text{UB}}(\delta)$ $|\cup_{P^{\text{ex}}\in\text{ex}\mathcal{P}}B_{P^{\text{ex}}}|$ $|\mathcal{H}|$ Realizability
$10$ $0.73777$ $0.76009$ $80$ $100$ Yes
$100$ $0.07549$ $0.07600$ $95$ $100$ Yes
$200$ $0.03780$ $0.03800$ $96$ $100$ Yes
$300$ $0.02520$ $0.02533$ $96$ $100$ Yes
$400$ $0.01892$ $0.01900$ $97$ $100$ Yes
$500$ $0.01514$ $0.01520$ $97$ $100$ Yes
Experiment 2: We conducted another synthetic experiment to show that the empirical risk for a given distribution is upper bounded by the traditional SLT bound of Corollary 4.3. This is a sanity check to see whether the environment we used in Experiment 1 is a valid one to check our results.
For the experiment, we selected a standard Gaussian distribution $\mathcal{N}(0,1)$ (mean 0, standard deviation = 1) to generate data. Similarly to Experiment 1, (i) the hypothesis space $\mathcal{H}$ is defined as a finite set of simple binary classifiers containing at least one realizable hypothesis, and (ii) we assume a zero-one loss function. The latter is used to evaluate the performance of the classifiers. For each run, we generated a training set and a test set from the standard Gaussian distribution. At training time, for each hypothesis $h$ in $\mathcal{H}$, we calculate the empirical risk on the training set using the zero-one loss and identify the hypothesis $\hat{h}$ that minimizes such risk (the empirical risk minimizer). At test time, we compute the empirical risk $L_{P}(\hat{h})$ of $\hat{h}$ on the test set as shown in Table 3.131313Here $P$ denotes the probability measure whose pdf is the standard Normal density.
We calculate the upper bound $\epsilon_{\text{UB}}(\delta)$ on the empirical risk based on Corollary 4.3, which is a function of the number of hypotheses in $\mathcal{H}$, the value of $\delta$, and the number $n$ of training samples. This is the classic SLT bound. The experiment is run $1000$ times with specified numbers of training and test samples, and a $\delta$ value of $0.05$ to check whether the condition (empirical risk on the test set is upper bounded by the theoretical bound) is satisfied in any of the $1000$ trials. The experimental results in Table 3 validate the classical SLT bound by repeatedly testing it on randomly generated data. The code for this experiment is also available upon request.
Table 3: Results of experimental evaluation. The hypothesis space is such that $|\mathcal{H}|=100$, and $\delta=0.05$.
Training Samples Test Samples $L_{P}(\hat{h})$ (Test) $\epsilon^{\star}(\delta)$ (Test) $\epsilon_{\text{UB}}(\delta)$ (Test)
$1000$ $500$ $0.000$ $0.01514$ $0.01520$
$1500$ $1000$ $0.000$ $0.00757$ $0.00760$
$2000$ $1500$ $0.000$ $0.00506$ $0.00506$
Appendix C A Simple Numerical Example for the $\epsilon$-Contamination Model of Section 3.1.1
Just like in Section 3.1.4, let $\Omega=\{\omega_{1},\omega_{2},\omega_{3}\}$, where $\omega_{j}=(x_{j},y_{j})$, $j\in\{1,2,3\}$. Suppose also that we observed four training samples $D_{1},\ldots,D_{4}$ and that we specified the likelihoods $\mathcal{L}_{1},\ldots,\mathcal{L}_{4}$ as in the following Table.
$\{\omega_{1}\}$$\{\omega_{2}\}$$\{\omega_{3}\}$
$\mathcal{L}_{1}$ $0.3$ $0.1$ $0.6$
$\mathcal{L}_{2}$ $0.4$ $0.2$ $0.4$
$\mathcal{L}_{3}$ $0.1$ $0.8$ $0.1$
$\mathcal{L}_{4}$ $0.15$ $0.7$ $0.15$
Then, suppose that $\epsilon_{1}=0.2$, $\epsilon_{2}=0.3$, $\epsilon_{3}=0.1$, and $\epsilon_{4}=0.25$, so that $\mathscr{L}_{1}=\{P:P=0.8\mathcal{L}_{1}+0.2Q\text{, }\forall Q\}$, $\mathscr{L}_{2}=\{P:P=0.7\mathcal{L}_{2}+0.3Q\text{, }\forall Q\}$, $\mathscr{L}_{3}=\{P:P=0.9\mathcal{L}_{3}+0.1Q\text{, }\forall Q\}$, and $\mathscr{L}_{4}=\{P:P=0.75\mathcal{L}_{1}+0.25Q\text{, }\forall Q\}$. By Wasserman and Kadane [1990, Example 3], we know that, if we put $\mathcal{P}=\text{Conv}(\cup_{i=1}^{4}\mathscr{L}_{i})$, the following holds
$\underline{P}(A)=\begin{cases}\min_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A),&\forall A\neq\Omega\\ 1,&\text{if }A=\Omega\end{cases}\quad\text{and}\quad\overline{P}(A)=\begin{cases}\max_{i\in\{1,\ldots,4\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)+\epsilon_{i},&\forall A\neq\emptyset\\ 0&\text{if }A=\emptyset\end{cases}.$
Simple calculations, then, give us the following values
$\underline{P}$ $\overline{P}$
$\{\omega_{1}\}$ $0.09$ $0.58$
$\{\omega_{2}\}$ $0.08$ $0.82$
$\{\omega_{3}\}$ $0.09$ $0.68$
$\{\omega_{1},\omega_{2}\}$ $0.32$ $0.91$
$\{\omega_{2},\omega_{3}\}$ $0.42$ $0.91$
$\{\omega_{1},\omega_{3}\}$ $0.18$ $0.92$
$\{\omega_{1},\omega_{2},\omega_{3}\}$ $1$ $1$
As we can see, in this example too the probability bounds imposed by the credal set are not too stringent, and in line with the evidence encapsulated in $\mathcal{L}_{1},\ldots,\mathcal{L}_{4}$. Hence, the assumption that $P^{\text{true}}\in\mathcal{P}$ is very plausible. For a visual representation of the credal set $\mathcal{P}=\text{Conv}(\cup_{i=1}^{4}\mathscr{L}_{i})$, see the yellow convex region in the next figure; it is very similar to the convex region in Figure 2. This is unsurprising, since the evidence used to derive the credal set in Section 3.1.4 is the same that we use to elicit $\mathcal{P}$ here.